### v1

# Creative Synthesis of Kinematic Mechanisms

Abstract
In this paper, we formulate the problem of kinematic synthesis for planar linkages as a cross-domain image generation task. We develop a planar linkages dataset using RGB image representations, covering a range of mechanisms: from simple types such as crank-rocker and crank-slider to more complex eight-bar linkages like Jansen’s mechanism. A shared-latent variational autoencoder (VAE) is employed to explore the potential of image generative models for synthesizing unseen motion curves and simulating novel kinematics. By encoding the drawing speed of trajectory points as color gradients, the same architecture also supports kinematic synthesis conditioned on both trajectory shape and velocity profiles. We validate our method on three datasets of increasing complexity: a standard four-bar linkage set, a mixed set of four-bar and crank-slider mechanisms, and a complex set including multi-loop mechanisms. Preliminary results demonstrate the effectiveness of image-based representations for generative mechanical design, showing that mechanisms with revolute and prismatic joints, and potentially cams and gears, can be represented and synthesized within a unified image generation framework. Code and dataset are available at: GitHub, Hugging Face.
1 Introduction
Kinematic synthesis is a long-standing problem in mechanical engineering. The objective is to design mechanisms that can generate a desired motion trajectory, using mechanical components such as links, sliders, gears, and cams. Within this broad area, the synthesis of planar linkages represents an important but still challenging subproblem. Classic examples such as the Watt’s and Stephenson’s linkages luo2018reconfigurable ; bai2016unified , as well as more complex designs like Jansen’s mechanism nansai2015jansen , illustrate how carefully designed linkages can produce sophisticated and purposeful motion. However, the synthesis problem remains challenging due to the complex and irregular nature of the design space. This complexity arises from the need to ensure a specific degree of freedom, manage redundant components, avoid motion singularities maloisel2021singularity , and account for multiple mechanisms that may produce the same trajectories zarkandi2021novel . These challenges span both the analysis perspective, simulating trajectories from given mechanisms, and synthesis perspective, generating mechanisms from motion curves. We demonstrate these symmetric processes in Figure 1.
Analytical approaches to kinematic synthesis date back to the 19th century, most notably with Burmester theory ceccarelli2006burmester , which provides closed-form geometric solutions for linkage design under discrete pose constraints. With the advent of computational methods, optimization-based techniques such as evolutionary algorithms lipson2006relaxation ; lipson2008evolutionary and interactive design tools like LinkEdit bacher2015linkedit have enabled the synthesis of more complex and non-intuitive mechanisms. More recently, the availability of large-scale datasets heyrani2022links ; nurizada2025dataset has enabled learning-based methods that train neural networks, including fully connected architectures, VAEs higgins2017beta ; deshpande2019computational ; nurizada2025path , and multi-modal encoders like CLIP radford2021learning ; nobari2024link , to predict mechanisms from motion examples. However, existing methods often rely on task-specific data structures such as graphs or joint coordinate lists, which restrict generalization across different types of mechanisms. In contrast, image-based representations provide a unified and scalable format that naturally encodes both motion and structure, enabling the same model architecture and training pipeline to be applied across a wide variety of kinematic structures.
In this work, we propose a cross domain generative framework that jointly models mechanical structures and their motion trajectories. Trained on a curated dataset of paired images of mechanisms and curves, the model learns a shared latent representation that bridges the two domains. Once trained, it enables both kinematic synthesis and analysis. Our contributions are as follows: 1). A new dataset of paired RGB images of planar mechanisms and their corresponding motion trajectories, covering diverse linkage types. 2). An end-to-end image generative model that learns a shared latent representation across the mechanism and trajectory domains. 3). Demonstration of cross domain synthesis and analysis, enabling high fidelity translation from trajectory to mechanism and from mechanism to trajectory.
Figure 1: We propose an image based generative framework for kinematic synthesis and analysis. The model supports both synthesis and analysis using the same architecture and representation. Trajectory speed is encoded using color gradients (left), while types of links and joints are encoded using predefined colors (right).
2 Related Work
Paper Dataset Curve Rep. Mechanism Rep. Method
Lipson lipson2008evolutionary (2008) Not fixed Coordinates Operator tree Genetic Programming
Vermeer et al.vermeer2018kinematic (2018) Not fixed Coordinates Operator tree Reinforcement Learning
Deshpande et al.deshpande2021image (2021) 4-links and 6-links Image Coordinates VAE+KNN
Pan et al.pan2023joint (2023) Not fixed Coordinates Graph Optimization
Fogelson et al.fogelson2023gcp (2023) Max 16-links Coordinates Graph GCN+Reinforcement Learning
Nobari et al.heyrani2022links ; nobari2024link (2024) Max 20-links Coordinates Graph CLIP+BFGS
Lee et al.lee2024deep (2024) 4-links Image Graph cGAN
Nurizada et al.nurizada2025path (2025) 4-links & sliders Image Coordinates $\beta$-VAE
Ours Max 12-links & sliders Image & Velocities Image or Video Shared-latent VAEs
Table 1: Summary of selected relevant kinematic synthesis papers.
Kinematic synthesis. Early work attempted to solve kinematic synthesis by interpolation. Simulation was used to generate a large database of mechanisms, and a neural network guided synthesis by interpolating a new mechanism from similar ones in the database vasiliu2001dimensional . Other approaches used a tree-based mechanism representation, relying on genetic programming lipson2006relaxation or reinforcement learning fogelson2023gcp to search the space of trees and find mechanisms with the desired properties.
Recent work has mostly focused on deep learning. A variety of generative models have been explored, including cGANs lee2024deep and VAEs nurizada2025path . Contrastive Learning has also been applied to enable rapid retrieval from massive mechanism databases nobari2024link . The rise of data driven approaches has further spurred the release of several high quality datasets  nurizada2025dataset ; heyrani2022links and even attempts at 3D synthesis cheng2022exact . Related to synthesis, mechanism design tools have also been explored. One example is mechanism editing, whereby a user can interact with an existing mechanism, and the solver allows for the user to add new properties while specifying properties that need to be preserved bacher2015linkedit . Another tool automated conversion from hand drawn sketches to digital representations nurizada2024transforming . Yet another presented an interactive design system for the creation of mechanical characters coros2013computational and linkage based characters thomaszewski2014computational , and multi-stable structures zhang2021computational .
Generative models have progressed from Variational Autoencoders (VAEs) kingma2013auto , which learn probabilistic latent representations, to Diffusion Models such as DDPMs ho2020denoising , which iteratively transform noise into data through denoising steps. Latent Diffusion Models (LDMs) rombach2022high perform diffusion in a compressed latent space, while Flow Matching lipman2022flow offers a continuous-time formulation that learns transformation between distributions. Cross-domain generative models aim to translate data from one domain to another while preserving semantic consistency. Applications includes image-to-image translation taigman2017unsupervised ; kim2017learning , few-shot domain adaptation ojha2021few , and cross-modal 3D synthesis long2024wonder3d . Multi-modal generative models extend this idea to learning joint representations across different modalities, such as video and audio ngiam2011multimodal . By jointly modeling multiple domains or modalities, these methods enable richer synthesis capabilities and bidirectional translation between diverse input types. Previous works have primarily focused on datasets such as SVHN to MNIST hoffman2018cycada or style translation richardson2021encoding between domains. In this work, we introduce kinematic synthesis as a new cross-domain generative modeling task, well-defined and application-relevant, and present a data generation pipeline with preliminary results on three datasets using the shared-latent VAEs.
3 Approach
Figure 2: Overview of our shared-latent VAE framework for cross-domain kinematic synthesis and analysis. Curve and mechanism images ($C$, $M$) are encoded into latent embeddings ($Z_{c}$, $Z_{m}$), which are aligned in a shared latent space. During training, both reconstruction ($\hat{C}$, $\hat{M}$) and cross-domain prediction ($\hat{C}_{m}$, $\hat{M}_{c}$) are supervised via back-propagation. At inference time, the model enables synthesis (from $C$ to $M^{\prime}$) and analysis (from $M$ to $C^{\prime}$) through feed-forward decoding. Following MAE he2022masked , we adopt an asymmetric ViT encoder–decoder design.
3.1 Shared-latent VAEs
Overview of our method shown in Figure 2. Let $C$ and $M$ denote the input curve and mechanism images. $E_{c}$, $E_{m}$ are the encoders, and $D_{c}$, $D_{m}$ are the decoders for the curve and mechanism branches. The latent codes are $Z_{c}=E_{c}(C)$ and $Z_{m}=E_{m}(M)$, sampled from a shared latent distribution $Z\sim\mathcal{N}(0,I)$. The reconstructions are $\hat{C}=D_{c}(Z_{c})$ and $\hat{M}=D_{m}(Z_{m})$, while the cross-domain predictions are $\hat{M}_{c}=D_{m}(Z_{c})$ and $\hat{C}_{m}=D_{c}(Z_{m})$. The training loss is defined as:
$\mathcal{L}_{\text{Shared-latent-VAEs}}=\$$\|C-\hat{C}\|^{2}+\|M-\hat{M}\|^{2}+\beta\cdot\left[\text{KL}(q(Z_{c}|C)\|p(Z))+\text{KL}(q(Z_{m}|M)\|p(Z))\right]$
$+\lambda\cdot\|Z_{c}-Z_{m}\|^{2}+\gamma\cdot(\|\hat{M}_{c}-M\|^{2}+\|\hat{C}_{m}-C\|^{2})$
The loss function consists of five components. The reconstruction loss measures the image reconstruction error for both the curve and mechanism domains, ensuring that each encoder-decoder pair can accurately reproduce its input. The KL divergence term regularizes the latent codes $Z_{c}$ and $Z_{m}$ to follow a standard normal distribution, promoting smoothness and structure in the latent space. The latent similarity loss encourages the curve and mechanism embeddings to be close, facilitating a shared latent representation across domains. Lastly, the cross-domain prediction loss ensures that the latent codes contain sufficient information to decode meaningful counterparts in the other domain, supporting both synthesis and analysis tasks.
3.2 Dataset
Figure 3: Examples of dataset construction. Left: Example mechanisms (4-bar mechanism and Jansen’s linkage) built with 2 and 5 triangle layers, with their corresponding sequences of link connections. Right: For each complexity level (T2–T5), two representative graphs from different isomorphism classes are shown, along with rendered mechanism examples from the dataset.
We construct a dataset of 1-DOF planar mechanisms by recursively applying triangle-based operators, starting from two fixed joints and one input joint, as shown in Figure 3. There are multiple operators to preserve the 1-DOF property while expanding the mechanism, such as T-operator and D-operator lipson2008evolutionary . Following recent work, we use the J-operator from the LINKS dataset heyrani2022links and the recurrent triangle solver introduced in LinkEdit bacher2015linkedit to build mechanisms stacking triangle layers. Our goal is to include historically significant mechanisms such as Watt’s, Stephenson’s, and Jansen’s, all of which share key properties: they begin with two fixed nodes, have one drawing node, and can be constructed with triangle layers. ***Sub-variants such as Watt-II and Stephenson-III, which involve three fixed nodes, are not included in the current version dataset. The dataset can be extended to include these cases by adjusting the filtering parameters.
To generate a compact yet expressive dataset spanning from simple four-bar linkages to complex mechanisms such as Jansen’s, we start with all possible combinations and then apply a set of structural filters. We begin with three initial nodes and recursively add one node at each layer by selecting two existing nodes as parents. This results in a total of $\prod_{i=3}^{k+2}\binom{i}{2}$ combinations for $k$ triangle layers, which corresponds to the first row in Table 2. For example, when $k=2$, the number of combinations is $\binom{3}{2}\cdot\binom{4}{2}=18$; when $k=5$, $\binom{3}{2}\cdot\binom{4}{2}\cdot\binom{5}{2}\cdot\binom{6}{2}\cdot\binom{7}{2}=56{,}700$. We only retain mechanisms that satisfy the following conditions: 1) a single drawing node, since mechanisms with multiple output points can be described by simpler substructures; 2) no redundant triangles, as adding a triangle onto an already rigid structure does not increase structural variety; 3) exactly two fixed nodes, to match the properties of the target mechanisms and avoid introducing additional grounded points; and 4) one representative graph per isomorphic class, removing any structure that differs only by the sequence of construction. As a result, we include 396 distinct linkage structures, along with an equal number of crank-slider-based structures obtained by changing the starting point of graph construction.
Filters | Triangle layers T0 T1 T2 T3 T4 T5 Total
Initial: All combinations 1 3 18 180 2700 56700 59602
Filter 1: One drawing node 0 1 5 31 257 2803 3097
Filter 2: Two fixed nodes 0 0 1 11 107 1227 1346
Filter 3: No redundant links 0 0 1 8 68 632 709
Filter 4: Isomorphic graphs 0 0 1 7 47 341 396
Table 2: Number of graphs retained after applying each structural filter, for the number of triangle layers from 0 (T0) to 5 (T5).
Color encoding is applied to both curve and mechanism images. For curves, the drawing-end speed is mapped to a color scale. For mechanisms, each component is assigned a predefined color: the base link is red, the input link is green, other links are blue, and joints are yellow. This encoding enables explicit validation when deriving parametric models. Inspired by the design of the Strandbeest’s nansai2015jansen walking leg, we observe that flattened contact points and steady horizontal motion of the output nodes promote stable forward movement and long locomotion overlap. Incorporating speed information into kinematic synthesis is both beneficial and easily integrated in RGB image representations.
4 Experiments
Figure 4: Qualitative results. Columns: (1) Curve to synthesized mechanism, compared with ground-truth mechanism; (2) Mechanism to calculated curve, shown beside the ground-truth curve; (3) Curve-to-mechanism-to-curve, compared with ground-truth curve. Rows list ViT-tiny and ViT-base dosovitskiy2020image models on three datasets.
We open-source the data generation code and provide three example datasets used to validate the proposed shared-latent VAE for image-based synthesis and analysis. Dataset-1 contains 100K four-bar mechanisms, including both crank–rocker and double–crank cases, referred to as T2. Dataset-2 contains 200K mixed four-bar mechanisms and crank–slider mechanisms, referred to as T2ST2. Dataset-3, denoted as T4, is a subset of our complete collection of distinct graphs, comprising 55 graphs up to T4, with 10K samples collected for each graph. We repeated the data collection process and used the resulting test set for all three experiments.
Models Dataset 1 (T2) PSNR $\uparrow$ Dataset 2 (T2 & ST2) PSNR $\uparrow$
Synthesis Analysis C2M2C Synthesis Analysis C2M2C
ViT-tiny 27.63 ± 2.28 26.59 ± 2.21 26.17 ± 2.36 29.48 ± 3.27 27.97 ± 2.60 28.01 ± 2.72
ResNet-18 28.58 ± 3.15 26.83 ± 2.61 27.03 ± 2.43 30.83 ± 4.54 28.20 ± 3.00 27.82 ± 3.34
ViT-base 28.54 ± 3.26 26.90 ± 2.43 27.18 ± 2.22 30.87 ± 4.76 28.43 ± 3.07 28.40 ± 3.20
Table 3: Baseline comparison. PSNR (mean ± std) for image accuracy in synthesis, analysis, and curve-to-mechanism-to-curve (C2M2C) tasks, on two datasets for three models dosovitskiy2020image ; he2016deep .
Ablations Dataset 1 (T2) PSNR $\uparrow$ Dataset 2 (T2 & ST2) PSNR $\uparrow$
Synthesis Analysis C2M2C Synthesis Analysis C2M2C
Black-white input 25.61 ± 3.10 23.42 ± 2.40 22.86 ± 2.35 27.38 ± 3.91 24.70 ± 2.83 24.22 ± 3.24
w/o analysis loss 26.43 ± 1.87 12.50 ± 0.38 12.49 ± 0.38 29.03 ± 3.20 12.52 ± 0.33 12.52 ± 0.33
w/o synthesis loss 12.41 ± 0.29 26.30 ± 1.89 18.55 ± 1.46 12.55 ± 0.30 27.01 ± 2.41 18.71 ± 1.28
w/o cross-domain loss 18.21 ± 0.96 19.22 ± 1.17 18.63 ± 1.73 18.61 ± 1.26 18.91 ± 1.57 19.03 ± 1.56
Complete loss (ViT-tiny) 27.63 ± 2.28 26.59 ± 2.21 26.17 ± 2.36 29.48 ± 3.27 27.97 ± 2.60 28.01 ± 2.72
Table 4: Ablation study. Effect of input color and loss function terms on PSNR (mean ± std) for synthesis, analysis, and C2M2C tasks.
Figure 5: Samples from the latent space, shown in a $3\times 3$ grid: from left to right: ViT-base with $\beta=10$ (curves, mechanisms) and CNN with $\beta=1$ (curves, mechanisms).
Qualitative results. We validate our method on three tasks: curve-to-mechanism synthesis, mechanism-to-curve analysis, and a two-stage generation where synthesis is followed by analysis. Figure 4 presents results for all three tasks across the three datasets. ViT-base generally produces more accurate outputs than ViT-tiny, particularly as the complexity of the mechanisms increases.
Quantitative results. Tables 3 and 4 report PSNR for all tasks. ViT-base achieves the highest scores across datasets compared with ResNet and ViT-tiny, especially in two-stage generation. Ablations show that removing synthesis or cross-domain losses significantly reduces accuracy, and color input outperforms black-and-white. The complete loss configuration yields the best overall performance.
5 Discussion
Limitations and future work. Figure 5 shows latent space samples for curves and mechanisms using ViT-base with $\beta=10$ and ResNet with $\beta=1$. The current sampling quality is limited, with blurred outputs and low-quality edges. This suggests that the VAE structure with high-capacity decoders may not have learned a well-structured normal distribution in the latent space.
In future work, large-scale datasets and more advanced generative models are encouraged for this task. In our data generation pipeline, we provide the optional recordings of joint coordinates and videos of one period’s mechanism motion, enabling future exploration of image-to-graph and image-to-video generation. Image-based representations for complex mechanism structures suffer from occlusion and ambiguity, whereas video representations can provide richer temporal and structural information.
Conclusion. In this paper, we introduced kinematic synthesis as a cross-domain generative modeling task and proposed a shared-latent VAE framework for translating between mechanisms and motion curves. We constructed a scalable dataset covering simple to complex planar mechanisms and demonstrated synthesis, analysis, and two-stage generation on multiple datasets. Preliminary results show that image-based representations enable a unified representation and end-to-end generative synthesis of kinematic mechanisms. This work lays the groundwork for applying generative modeling into mechanism design, with potential applications in robotics.
Appendix A Additional Results: Real-World Validation
Figure 6: Examples of mechanism–curve pairs and real-world validation. Top left: a dataset example showing a mechanism and its corresponding drawing curve. Top right: synthesis (curve-to-mechanism) and analysis (mechanism-to-curve) results compared with ground truth, using the ViT-Tiny model. Bottom: a 3D-printed mechanism driven by a servo motor and the traced drawing trajectory of its end-effector.